# RLHF and IIA: Perverse Incentives

### 02 Dec 2023 v1

@@ 4.2 Direct Preference Optimization
For a parameterized language model $\widehat{\pi}_{\theta}$, minimizing $\mathcal{L}_{\mathrm{DPO}}$ produces a fine-tuned language model $\widehat{\pi}_{\widehat{\theta}}$.
As we will see, RLPO and DPO both can give rise to egregious behavior when queries include more than two alternatives. Recent work proposes a family of RLHF algorithms that unifies and extends RLPO and DPO [3]. Other so-called $\Psi$PO algorithms suffer in the same manner when queries include more than two alternatives.
@@ 4.3 Inclusive Learning
RLPO and DPO maximize reward. This tends toward monolithic language models that serve a majority of the population. Inclusive learning (IL) offers an alternative approach to RLHF, aiming to produce inclusive language models that reflect the diversity of preferences across the population [14].
IL produces a language model $\pi$ that simultaneously serves as a reward model. In particular, reward is taken to be the log-probability $\ln P_{\pi}(x)$ assigned to a message. Minimizing cross-entropy loss while regularizing toward the base policy $\overline{\pi}$ gives rise to an inclusive loss function
@@ 5.2 Architectures
Under this assumption, each reward function $r_{\psi}$ is parameterized by two scalars – $\psi_{1}$ and $\psi_{2}$ – which express estimates of rewards enjoyed by individuals who receive their desired type of message.
Our next assumption pertains to the policy architecture.
@@ 5.4 RLPO and DPO
These equations imply that $\widehat{\psi}_{2}-\widehat{\psi}_{1}=\ln(4/3)>0$. Hence, the estimated reward $\widehat{r}_{\widehat{\psi}}(x)$ is maximized, as RLPO aims to do, by generating messages $x\in\mathcal{M}_{2}$. DPO is designed to approximate RLPO and therefore leads to similar behavior.
In our thought experiment, the fact that each choice set had twice as many more desirable ($\mathcal{M}_{1}$) than the less desirable ($\mathcal{M}_{2}$) messages biased reward estimates in favor of $\mathcal{M}_{2}$. This tendency is expressed in the above propositions though the requirement that $F(P_{\overline{\pi}}(\mathcal{M}_{1}))>p_{*}(1)$, with $F(\cdot)$ defined in Propositions 4 and 5. The baseline policy $\overline{\pi}$ is used to generate choice sets, and perhaps surprisingly, the fact that it biases samples toward desirable messages leads to undesirable outcomes.
The root cause is that $\widehat{r}_{\widehat{\psi}}$ assumes IIA while the process generating preferences does not. In our thought experiment, for example, where $\mathbb{P}(Y=1|\mathcal{Y})=\mathbb{P}(Y=2|\mathcal{Y})=0.3$ and $\mathbb{P}(Y=3|\mathcal{Y})=0.4$, if the choice sets were instead to only contain two alternatives $\mathcal{Y}=\{y_{2},y_{3}\}$ of different types, then we would have $\mathbb{P}(Y=2|\mathcal{Y})=0.6$ and $\mathbb{P}(Y=3|\mathcal{Y})=0.4$, giving rise to different equations
@@ 5.5 Inclusive Learning
Hence, $P_{\widehat{\pi}_{\widehat{\theta}}}(\mathcal{M}_{1})=15/(15+|\mathcal{M}_{2}|)$. It follows that, as $\mathcal{M}_{2}$ grows, $P_{\widehat{\pi}_{\widehat{\theta}}}(\mathcal{M}_{1})$ vanishes.
While the number of messages in $\mathcal{M}_{2}$ does not influence how a human would choose between two messages, (17) implies that it impacts the probabilities $P_{\widehat{\pi}_{\widehat{\theta}}}(\mathcal{M}_{1})$ and $P_{\widehat{\pi}_{\widehat{\theta}}}(\mathcal{M}_{2})$ of generating more or less liked messages. The more equivalent messages there are in $\mathcal{M}_{2}$, the more likely the resulting fine-tuned language model is to produce less-liked messages. This is again due to the fact that the model underlying IL satisfies IIA while the preference data generating process does not.
@@
This observation suggests that using IL to fine-tune real language models will induce a bias toward generating longer messages. This is because the number of ways to express information scales quickly with the length of message required to express that information. To see why, suppose there are two ideas that might be expressed as responses to a prompt, and that the first requires a about ten words, while the second requires about one thousand words. Suppose that $60\%$ of the population would prefer the first idea while $40\%$ would prefer the second. Because there are so many more thousand word than ten word statements, there are likely to be many more roughly equivalent expressions of the second idea than the first. Since IL biases generation toward messages with many equivalent alternatives, the fine-tuned language model would tend to express the second idea.
@@ 6 Closing Remarks
The essential role of RLHF in the success of generative AI has attracted tremendous effort and resources to the subject. It is worthwhile to think carefully about foundations of this methodology to inform algorithmic innovation and ultimately produce reliable AIs. What we have presented, which points out how the IIA assumption adopted by models underlying current algorithms can give rise to egregious behavior, offers one of what will hopefully be many steps in this direction.
As we discussed, in some plausible contexts, IIA incentivizes generation of verbose messages when concise ones are preferred. Recent work studies how RLHF is giving rise to messages longer than desired [11]. While that work uncovers other culprits of this phenomenon, it may be worth exploring the relation to ideas we have presented.
The design of RLHF algorithms that mitigate perverse incentives we have identified remains a subject for future research. Trade-offs between query formats and how feedback is processed also deserve further study. For example, while RLPO can fail when queries present more than two alternatives, one can heuristically convert a feedback from such a query to multiple pairwise choices. This could offer more robust results, though there may be other flaws to this approach.
@@ A.1 RLPO
In this subsection, we provide the proof for Proposition 4. The proof can be split into three steps. First, we prove a useful lemma that shows that, if the proportion of sampled category 1 messages is larger than the proportion of preferred category 1 messages in the data, then RLPO will produce higher rewards for category 2 messages.
Step 1: RLPO produces higher rewards for category 2 messages if data is biased toward more category 1 messages.
We start by defining some shorthand notations. To simplify the exposition, we define $N=|\mathcal{D}|$ and $K=|\mathcal{Y}|$. For each $n=1,2,\ldots,N$, we use $\left(\mathcal{Y}^{(n)},y^{(n)}\right)$ to denote the $n$-th datum in $\mathcal{D}$. Moreover, for each category $\tau=1,2$, we define $q_{\tau}$ as the proportion of data in $\mathcal{D}$ where messages of category $\tau$ is chosen as the preferred message:
$q_{\tau}=\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}\left(y^{(n)}\in\mathcal{M}_{\tau}\right)\quad\forall\tau=1,2.$
Moreover, for each $n=1,2,\ldots,N$, we use $k_{\tau}^{(n)}$ to denote the number of messages in $\mathcal{Y}^{(n)}$ of category $\tau$. That is
$k_{\tau}^{(n)}=\sum_{y\in\mathcal{Y}^{(n)}}\mathbbm{1}(y\in\mathcal{M}_{\tau})\qquad\forall n=1,2,\ldots,N\quad\textrm{ and }\quad\forall\tau=1,2$
Note that $\frac{1}{NK}\sum_{n=1}^{N}k_{1}^{(n)}$ is the proportion of category 1 messages in data $\mathcal{D}$. We say $\mathcal{D}$ is biased toward category 1 messages if the proportion of category 1 messages in $\mathcal{D}$ is larger than $q_{1}$, the proportion of data that category 1 messages are chosen. The following lemma formalizes the result of Step 1.
@@
If $0<q_{1}\leq 1$ and
$\frac{1}{NK}\sum_{n=1}^{N}k_{1}^{(n)}>q_{1},$
then the minimizer $\widehat{r}$ of $\mathcal{L}_{\mathrm{reward}}$ exists and satisfies
$\widehat{r}(x^{\prime})-\widehat{r}(x)>1-\left[1+K^{2}\cdot\left(\frac{\sum_{n=1}^{N}k_{1}^{(n)}k_{2}^{(n)}}{N}\right)^{-1}\cdot\left(\frac{1}{NK}\sum_{n=1}^{N}k_{1}^{(n)}-q_{1}\right)\right]^{-1}\quad\text{for all }x\in\mathcal{M}_{1},x^{\prime}\in\mathcal{M}_{2}.$
@@
For simplicity, for $x\in\mathcal{M}_{1}$ and $x^{\prime}\in\mathcal{M}_{2}$, we use $r_{1}$ and $r_{2}$ as shorthands for $r(x)$ and $r(x^{\prime})$, and $\widehat{r}_{1}$ and $\widehat{r}_{2}$ for $\widehat{r}(x)$ and $\widehat{r}(x^{\prime})$. Here the loss function can be written as
@@
$\mathcal{L}_{\mathrm{reward}}(r)$ $=-\sum_{(\mathcal{Y},y)\in\mathcal{D}}\ln\frac{e^{r(y)}}{\sum_{y^{\prime}\in\mathcal{Y}}e^{r(y^{\prime})}}$
$=-r_{1}Nq_{1}-r_{2}N(1-q_{1})+\sum_{n=1}^{N}\ln\left(k_{1}^{(n)}e^{r_{1}}+k_{2}^{(n)}e^{r_{2}}\right).$
@@
Note that this loss function is strictly convex and thus a local minima is also a global minima. Examining the partial derivatives, we can see that $\widehat{r}_{1}$ and $\widehat{r}_{2}$ satisfy
$Nq_{1}=\sum_{n=1}^{N}\frac{k_{1}^{(n)}e^{\widehat{r}_{1}}}{k_{1}^{(n)}e^{\widehat{r}_{1}}+k_{2}^{(n)}e^{\widehat{r}_{2}}}.$
We let $\zeta=e^{\widehat{r}_{2}-\widehat{r}_{1}}$, then
$f(\zeta)=\sum_{n=1}^{N}\frac{k_{1}^{(n)}}{k_{1}^{(n)}+k_{2}^{(n)}\zeta}=Nq_{1}.$
Note that $f(\zeta)$ is strictly decreasing in $\zeta$, and
$f(1)=\sum_{n=1}^{N}\frac{k_{1}^{(n)}}{k_{1}^{(n)}+k_{2}^{(n)}}=\frac{1}{K}\sum_{n=1}^{N}k_{1}^{(n)},$
@@
where we use the fact that $k_{1}^{(n)}+k_{2}^{(n)}=K$ for each $n$. Clearly, $f(\zeta)\to 0$ as $\zeta\to\infty$. Since $Nq_{1}>0$, there exists a unique $\zeta$ such that $f(\zeta)=Nq_{1}$. Further, we also note that $f(\zeta)$ is convex in $\zeta$, and
$f^{\prime}(1)=-\sum_{n=1}^{N}\frac{k_{1}^{(n)}k_{2}^{(n)}}{\Big{(}k_{1}^{(n)}+k_{2}^{(n)}\Big{)}^{2}}=-\frac{1}{K^{2}}\sum_{n=1}^{N}k_{1}^{(n)}k_{2}^{(n)}.$
Hence, for all $x>1$,
$Nq_{1}=f(\zeta)>f(1)+f^{\prime}(1)(\zeta-1)=\frac{1}{K}\sum_{n=1}^{N}k_{1}^{(n)}-(\zeta-1)\cdot\frac{1}{K^{2}}\sum_{n=1}^{N}k_{1}^{(n)}k_{2}^{(n)}.$ (18)
This leads to
@@
$\zeta-1$ $>$ $K^{2}\left(\sum_{n=1}^{N}k_{1}^{(n)}k_{2}^{(n)}\right)^{-1}\cdot\left(\frac{1}{K}\sum_{n=1}^{N}k_{1}^{(n)}-Nq_{1}\right)$ (19)
$=$ $K^{2}\left(\frac{\sum_{n=1}^{N}k_{1}^{(n)}k_{2}^{(n)}}{N}\right)^{-1}\cdot\left(\frac{1}{NK}\sum_{n=1}^{N}k_{1}^{(n)}-q_{1}\right).$
@@
And the result follows from combining (19) and the fact that $\ln(1+\alpha)\geq\frac{\alpha}{1+\alpha}=1-\frac{1}{1+\alpha}$ for all $\alpha>-1$. ∎
The second step is to show that, under Assumption 1, data is indeed biased toward more category 1 messages.
Step 2: With high probability, data is biased toward more category 1 messages.
We define
@@
as the subset of $\mathcal{D}$ that contains both categories of messages. The first lemma below shows that most of the data contains both categories of messages with high probability.
@@
Under Assumption 1, if $P_{\overline{\pi}}(\mathcal{M}_{1})\in(0,1)$, $K\geq 2$, then there exists a constant $\gamma=\gamma(K)\in(0,1)$ such that for all $N>0$, $\mathbb{P}(|\mathcal{I}|\leq N(1-\gamma))\leq e^{-\frac{N(1-\gamma)\gamma^{2}}{8}}$.
@@
Under Assumption 1, for each $(\mathcal{Y},y)\in\mathcal{D}$, the membership $\tau$ of each message $y^{\prime}\in\mathcal{Y}$ follows $\mathrm{Bernoulli}(P_{\overline{\pi}}(\mathcal{M}_{1}))$. Let $s_{K}=1-P_{\overline{\pi}}(\mathcal{M}_{1})^{K}-P_{\overline{\pi}}(\mathcal{M}_{2})^{K}<1$. Then $s_{K}$ equals the probability of sampling both categories of messages in each datum. Then, $|\mathcal{I}|\sim\mathrm{Binomial}(N,s_{K})$. By Chernoff’s bound,
$\mathbb{P}\left(|\mathcal{I}|\leq s_{K}\cdot Ns_{K}\right)\leq\exp\left(-\frac{Ns_{K}(1-s_{K})^{2}}{2}\right).$
Letting $\gamma=1-s_{K}^{2}\in(0,1)$, we have $s_{K}=\sqrt{1-\gamma}$ and $\gamma=(1-s_{K})(1+s_{K})<2(1-s_{K})$. Thus, $s_{K}(1-s_{K})^{2}>\sqrt{1-\gamma}(\gamma/2)^{2}>(1-\gamma)\gamma^{2}/4$. Therefore,
$\mathbb{P}\left(|\mathcal{I}|\leq N(1-\gamma)\right)\leq\exp\left(-\frac{N(1-\gamma)\gamma^{2}}{8}\right).$ (20)
The following lemma shows that when data is biased, the resulting proportion of sampled category 1 messages is greater than the proportion of chosen category 1 data with high probability.

### 21 Dec 2023 v2

@@ 4.2 Direct Preference Optimization
For a parameterized language model $\widehat{\pi}_{\theta}$, minimizing $\mathcal{L}_{\mathrm{DPO}}$ produces a fine-tuned language model $\widehat{\pi}_{\widehat{\theta}}$. As we will see, RLPO and DPO both can give rise to egregious behavior when queries include more than two alternatives. Recent work proposes a family of RLHF algorithms that unifies and extends RLPO and DPO [3]. Other so-called $\Psi$PO algorithms suffer in the same manner when queries include more than two alternatives.
@@ 4.3 Inclusive Learning
RLPO and DPO maximize reward. This tends toward monolithic language models that serve a majority of the population. Inclusive learning (IL) offers an alternative approach to RLHF, aiming to produce inclusive language models that reflect the diversity of preferences across the population [16]. IL produces a language model $\pi$ that simultaneously serves as a reward model. In particular, reward is taken to be the log-probability $\ln P_{\pi}(x)$ assigned to a message. Minimizing cross-entropy loss while regularizing toward the base policy $\overline{\pi}$ gives rise to an inclusive loss function
@@ 5.2 Architectures
Under this assumption, each reward function $r_{\psi}$ is parameterized by two scalars – $\psi_{1}$ and $\psi_{2}$ – which express estimates of rewards enjoyed by individuals who receive their desired type of message. Our next assumption pertains to the policy architecture.
@@ 5.4 RLPO and DPO
These equations imply that $\widehat{\psi}_{2}-\widehat{\psi}_{1}=\ln(4/3)>0$. Hence, the estimated reward $\widehat{r}_{\widehat{\psi}}(x)$ is maximized, as RLPO aims to do, by generating messages $x\in\mathcal{M}_{2}$. DPO is designed to approximate RLPO and therefore leads to similar behavior. In our thought experiment, the fact that each choice set had twice as many more desirable ($\mathcal{M}_{1}$) than the less desirable ($\mathcal{M}_{2}$) messages biased reward estimates in favor of $\mathcal{M}_{2}$. This tendency is expressed in the above propositions though the requirement that $F(P_{\overline{\pi}}(\mathcal{M}_{1}))>p_{*}(1)$, with $F(\cdot)$ defined in Propositions 4 and 5. The baseline policy $\overline{\pi}$ is used to generate choice sets, and perhaps surprisingly, the fact that it biases samples toward desirable messages leads to undesirable outcomes. The root cause is that $\widehat{r}_{\widehat{\psi}}$ assumes IIA while the process generating preferences does not. In our thought experiment, for example, where $\mathbb{P}(Y=1|\mathcal{Y})=\mathbb{P}(Y=2|\mathcal{Y})=0.3$ and $\mathbb{P}(Y=3|\mathcal{Y})=0.4$, if the choice sets were instead to only contain two alternatives $\mathcal{Y}=\{y_{2},y_{3}\}$ of different types, then we would have $\mathbb{P}(Y=2|\mathcal{Y})=0.6$ and $\mathbb{P}(Y=3|\mathcal{Y})=0.4$, giving rise to different equations
@@ 5.5 Inclusive Learning
Hence, $P_{\widehat{\pi}_{\widehat{\theta}}}(\mathcal{M}_{1})=15/(15+|\mathcal{M}_{2}|)$. It follows that, as $\mathcal{M}_{2}$ grows, $P_{\widehat{\pi}_{\widehat{\theta}}}(\mathcal{M}_{1})$ vanishes. While the number of messages in $\mathcal{M}_{2}$ does not influence how a human would choose between two messages, (17) implies that it impacts the probabilities $P_{\widehat{\pi}_{\widehat{\theta}}}(\mathcal{M}_{1})$ and $P_{\widehat{\pi}_{\widehat{\theta}}}(\mathcal{M}_{2})$ of generating more or less liked messages. The more equivalent messages there are in $\mathcal{M}_{2}$, the more likely the resulting fine-tuned language model is to produce less-liked messages. This is again due to the fact that the model underlying IL satisfies IIA while the preference data generating process does not. This observation suggests that using IL to fine-tune real language models will induce a bias toward generating longer messages. This is because the number of ways to express information scales quickly with the length of message required to express that information. To see why, suppose there are two ideas that might be expressed as responses to a prompt, and that the first requires a about ten words, while the second requires about one thousand words. Suppose that $60\%$ of the population would prefer the first idea while $40\%$ would prefer the second. Because there are so many more thousand word than ten word statements, there are likely to be many more roughly equivalent expressions of the second idea than the first. Since IL biases generation toward messages with many equivalent alternatives, the fine-tuned language model would tend to express the second idea.
6 Empirical Study
We next demonstrate that the sort of egregious behavior we have identified manifests when learning reward models of practical scale from realistic datasets. In particular, we fit reward models that build on the PaLM 2 [5] XXS language model to data generated using GPT-3.5 and GPT-4 [10]. Each query includes reponses generated by GPT-3.5 and GPT-4. We additionally use GPT-4 to simulate choices made by human annotators. Our results establish that, when training queries each include a pair of responses, the reward model correctly learns to assign higher scores to GPT-4 responses. However, when training queries each include four responses, the reward model erroneously assigns favorable scores to GPT-3.5 responses. It is striking that a seemingly innocuous change to the training query format gives rise to such egregious behavior. The results are summarized in Figure 4. Before discussing them in detail, we will describe the datasets, the reward model architecture, and the training algorithm.
(a) Test accuracy of a standard reward model trained on preference data with 2 or 4 responses per query. The dashed line indicates a purely random baseline.
(b) Percentage of test data where the reward model assigns a higher score to GPT-4’s response.
Figure 4: Standard reward model training can lead to egregious outcomes when training data involves more than two responses per query.
6.1 Preference Datasets
For simplicity, we consider only a single prompt. While other prompts lead to similar results, we arbitrarily choose to present results produced by the following prompt, based on a popular movie this year:
prompt Did Oppenheimer win a Nobel Prize?
We sample candidate responses using GPT-3.5 and GPT-4 with the temperature parameter of each language model set to the default value of one. Here are representative responses:
GPT-3.5 response No, Oppenheimer did not win the Nobel Prize. GPT-4 response No, Robert Oppenheimer, often called the “father of the atomic bomb” for his role in the Manhattan Project, did not win a Nobel Prize.
GPT-3.5 tends to produce concise responses relative to the more informative ones from GPT-4. From each of these two language models, we sample 100 responses for training and validation, and 100 responses for testing. We subsample from these responses when constructing each training or validation query, or each test sample. We produce two training sets that differ in the number of alternative responses per query. The first training set has 2 responses per query, one from GPT-3.5 and one from GPT-4. The second training set has 4 responses per query, one from GPT-3.5 and three from GPT-4. We generate 800 queries for each training set. Additionally, we generate 200 queries for each validation set. These are used in tuning hyperparameters to optimize validation accuracy. We produce a single test dataset made up of 1000 queries, each with one response from GPT-3.5 and one from GPT-4. To simulate human annotator choices, we prompt GPT-4 to select its favorite response among a set. We use two types of prompts, which express preference for concise and informative responses respectively:
concise choice prompt Suppose that you are looking for a concise answer to the following question. Which response below do you like the best… informative choice prompt Suppose that you are trying to learn more about the following question. Which response below do you like the best…
We find that prompting for a concise choice typically favors GPT-3.5, while prompting for an informative choice typically favors GPT-4. When simulating a human annotator, we sample one of these two types of choice prompts according to probabilities 0.3 and 0.7, respectively. Using these choice prompts, we find that regardless of the number of responses per query, our simulated human annotators select responses from GPT-4 approximately 70% of the time.
6.2 Reward Model Architecture and Training Algorithm
Our reward model builds on the PaLM 2 [5] XXS language model. In particular, given a context formed by concatenating a prompt and a response, we take the final embedding produced by the base language model and apply a linear layer to obtain reward. We train the linear layer, as well as the base language model, to minimize cross-entropy loss on preference data. For each training dataset, we train the reward model over 150 gradient steps using the Adafactor optimizer [13] with a learning rate of 1e-4 and batch size of 16. The hyperparameters are tuned to optimize validation accuracy. We average results over 3 seeds.
6.3 Results
Figure 3(a) plots the test accuracy of the reward model trained on queries with 2 or 4 responses. When each training query includes only a pair of responses, the reward model correctly selects the preferred response for around 70% of test samples. However, with four responses per training query, the test accuracy drops below 50%. In other words, choice predictions generated by the reward model fare worse than random guessing. Figure 3(b) provides more insight into how learned reward models score responses. Since the majority of simulated annotators favor GPT-4, we would expect that the learned reward model assigns higher scores to GPT-4 responses. We see in Figure 3(b) that, when trained on pairwise comparisons, the reward model almost always assigns higher scores to GPT-4 responses. However, when trained on queries each with 4 responses, the reward model tends to erroneously assign higher scores to GPT-3.5 responses. These results highlight a failing of the standard reward learning approach. The culprit is the IIA assumption made in reward learning, which is violated by the data as we will now explain. Intuitively, each simulated human annotator prefers either concise or informative responses but is relatively indifferent between alternatives in each of these two categories. Consider a simulated annotator who, when presented with one concise and one informative response, selects the former with probability 3/5. Under the IIA assumption, increasing the number of informative alternatives from one to three would reduce the probability of choosing the concise response to 1/3. However, if the annotator is indifferent between the three informative responses, the probability of choosing the concise response should remain 3/5, regardless of whether there are one or three informative alternatives.
@@ 7 Closing Remarks
The essential role of RLHF in the success of generative AI has attracted tremendous effort and resources to the subject. It is worthwhile to think carefully about foundations of this methodology to inform algorithmic innovation and ultimately produce reliable AIs. What we have presented, which points out how the IIA assumption adopted by models underlying current algorithms can give rise to egregious behavior, offers one of what will hopefully be many steps in this direction. The design of RLHF algorithms that mitigate perverse incentives we have identified remains a subject for future research. Trade-offs between query formats and how feedback is processed also deserve further study. For example, while RLPO can fail when queries present more than two alternatives, one can heuristically convert a feedback from such a query to multiple pairwise choices. This could offer more robust results, though there may be other flaws to this approach.
@@ A.1 RLPO
In this subsection, we provide the proof for Proposition 4. We start by defining two key quantities for messages of category $1$ in $\mathcal{D}$. First, we let
$\rho_{\mathrm{chosen}}=\frac{1}{|\mathcal{D}|}\sum_{(\mathcal{Y},y)\in\mathcal{D}}\mathbbm{1}\left(y\in\mathcal{M}_{1}\right)$
denote the fraction of data in $\mathcal{D}$ where the chosen message is in $\mathcal{M}_{1}$. Second, we let
$\rho_{\mathrm{data}}=\frac{1}{|\mathcal{D}|}\sum_{(\mathcal{Y},y)\in\mathcal{D}}\frac{|\mathcal{Y}\cap\mathcal{M}_{1}|}{|\mathcal{Y}|}$
denote the fraction of messages in $\mathcal{D}$ that belong to $\mathcal{M}_{1}$.
@@
For all strictly convex functions $f:\Re\to\Re$ such that $f^{\prime}(1)<0$, if $f(\zeta)=0$ for some $\zeta\in\Re$, then $\zeta>1-\frac{f(1)}{f^{\prime}(1)}$.
@@
Since $f$ is strictly convex, $f(\zeta)>f(1)+f^{\prime}(1)(\zeta-1)$. Since $f^{\prime}(1)<0$, this implies $\zeta-1>-\frac{f(1)}{f^{\prime}(1)}$. ∎
Lemma 8.
Under Assumption 2, if $\rho_{\mathrm{data}}>\rho_{\mathrm{chosen}}>0$, then the minimizer $\widehat{\psi}$ of $\mathcal{L}_{\mathrm{reward}}(\widehat{r}_{\psi})$ exists and satisfies
$\widehat{\psi}_{2}-\widehat{\psi}_{1}>\frac{\rho_{\mathrm{data}}-\rho_{\mathrm{chosen}}}{1+\rho_{\mathrm{data}}-\rho_{\mathrm{chosen}}}.$
Proof.
Since $\mathcal{L}_{\mathrm{reward}}$ is invariant under translations of $\psi$, without loss of generality, set $\psi_{1}=0$.
@@
$\mathcal{L}_{\mathrm{reward}}(\widehat{r}_{\psi})$ $=-\sum_{(\mathcal{Y},y)\in\mathcal{D}}\ln\frac{e^{\widehat{r}_{\psi}(y)}}{\sum_{y^{\prime}\in\mathcal{Y}}e^{\widehat{r}_{\psi}(y^{\prime})}}=-\psi_{2}|\mathcal{D}|(1-\rho_{\mathrm{chosen}})+\sum_{(\mathcal{Y},y)\in\mathcal{D}}\ln\left(|\mathcal{Y}\cap\mathcal{M}_{1}|+|\mathcal{Y}\cap\mathcal{M}_{2}|e^{\psi_{2}}\right).$
@@
Note that this loss function is strictly convex and thus a local minima is also a global minima. Examining the derivative, we see that $\widehat{\psi}_{2}$, if exists, satisfies
$\sum_{(\mathcal{Y},y)\in\mathcal{D}}\frac{|\mathcal{Y}\cap\mathcal{M}_{1}|}{|\mathcal{Y}\cap\mathcal{M}_{1}|+|\mathcal{Y}\cap\mathcal{M}_{2}|e^{\widehat{\psi}_{2}}}-|\mathcal{D}|\rho_{\mathrm{chosen}}=0.$
With a change of variable, let $x=e^{\psi_{2}}$ and
$f(x)=\sum_{(\mathcal{Y},y)\in\mathcal{D}}\frac{|\mathcal{Y}\cap\mathcal{M}_{1}|}{|\mathcal{Y}\cap\mathcal{M}_{1}|+|\mathcal{Y}\cap\mathcal{M}_{2}|x}-|\mathcal{D}|\rho_{\mathrm{chosen}}.$
@@
Clearly, $f(x)$ is strictly decreasing on $\Re_{+}$ and $f(x)\to 0$ as $x\to\infty$. Since $\rho_{\mathrm{chosen}}>0$, there exists a unique $\zeta>0$ such that $f(\zeta)=0$. Thus, $\widehat{\psi}_{2}=\ln\zeta$ exists. Further, we note that $f(x)$ is strictly convex and $f^{\prime}(1)<0$. By Lemma 7,
@@
$\zeta>1-\frac{f(1)}{f^{\prime}(1)}=1+|\mathcal{D}|(\rho_{\mathrm{data}}-\rho_{\mathrm{chosen}})\cdot|\mathcal{Y}|^{2}\left(\sum_{(\mathcal{Y},y)\in\mathcal{D}}|\mathcal{Y}\cap\mathcal{M}_{1}|\cdot|\mathcal{Y}\cap\mathcal{M}_{2}|\right)^{-1}>1+\rho_{\mathrm{data}}-\rho_{\mathrm{chosen}}.$ (18)
@@
The result follows from combining Equation (18) and the fact that $\ln(1+\alpha)\geq\frac{\alpha}{1+\alpha}$ for all $\alpha>-1$. ∎
For all $\eta>0$, define the high probability event
$\mathcal{E}_{\eta}(\mathcal{D})=\left\{\rho_{\mathrm{data}}-\rho_{\mathrm{chosen}}>\eta\right\}\bigcap\left\{\rho_{\mathrm{chosen}}>0\right\}.$ (19)
Lemma 9.
Under Assumptions 2 and 3, if there exists $\eta>0$ such that $\mathbb{P}(\mathcal{E}_{\eta}(\mathcal{D}))\to 1$ as $|\mathcal{D}|\to\infty$, then as $|\mathcal{D}|\to\infty$,
[…]
Proof.
By Lemma 8, under event $\mathcal{E}_{\eta}(\mathcal{D})$, $\widehat{\psi}=\operatorname*{arg\,min}_{\psi}\mathcal{L}_{\mathrm{reward}}(\widehat{r}_{\psi})$ and $\widehat{\psi}_{2}-\widehat{\psi}_{1}>\frac{\rho_{\mathrm{data}}-\rho_{\mathrm{chosen}}}{1+\rho_{\mathrm{data}}-\rho_{\mathrm{chosen}}}>\frac{\eta}{1+\eta}$. When $\beta=0$, $\widehat{\pi}_{\theta}$ maximizes reward. Thus, under event $\mathcal{E}_{\eta}(\mathcal{D})$, $P_{\widehat{\pi}_{\widehat{\theta}}}(\mathcal{M}_{1})=0$ and $P_{\widehat{\pi}_{\widehat{\theta}}}(\mathcal{M}_{2})=1$. When $\beta>0$, $\widehat{\pi}_{\theta}$ satisfies
$P_{\widehat{\pi}_{\theta}}(y)=\frac{P_{\overline{\pi}}(y)e^{\widehat{r}_{\widehat{\psi}}(y)|\mathcal{D}|/\beta}}{\sum_{y^{\prime}\in\mathcal{M}}P_{\overline{\pi}}(y^{\prime})e^{\widehat{r}_{\widehat{\psi}}(y^{\prime})|\mathcal{D}|/\beta}}.$
Under event $\mathcal{E}_{\eta}(\mathcal{D})$,
$\frac{P_{\widehat{\pi}_{\theta}}(\mathcal{M}_{1})}{P_{\widehat{\pi}_{\theta}}(\mathcal{M}_{2})}=\frac{p_{1}}{p_{2}}\cdot e^{\left(\widehat{\psi}_{1}-\widehat{\psi}_{2}\right)|\mathcal{D}|/\beta}<e^{-\frac{\eta}{1+\eta}\cdot\frac{|\mathcal{D}|}{\beta}}.$
For all $\epsilon>0$, there exists $N>0$ such that $e^{-\frac{\eta}{1+\eta}\cdot\frac{|\mathcal{D}|}{\beta}}<\epsilon$ for all $|\mathcal{D}|>N$. In turn, for all $\beta\geq 0$ and $\epsilon>0$, there exists $N>0$ such that for all $|\mathcal{D}|>N$,
$\mathbb{P}\left(\frac{P_{\widehat{\pi}_{\theta}}(\mathcal{M}_{1})}{P_{\widehat{\pi}_{\theta}}(\mathcal{M}_{2})}\geq\epsilon\right)\leq\mathbb{P}\left(\mathcal{E}_{\eta}(\mathcal{D})^{c}\right).$
The proof follows from the fact that $\mathbb{P}(\mathcal{E}_{\eta}(\mathcal{D})^{c})\to 0$ as $|\mathcal{D}|\to\infty$ and $P_{\widehat{\pi}_{\theta}}(\mathcal{M}_{1})+P_{\widehat{\pi}_{\theta}}(\mathcal{M}_{2})=1$. ∎
Finally, we prove that under the dichotomy data assumption, there exists $\eta>0$ such that the event $\mathcal{E}_{\eta}(\mathcal{D})$ holds with high probability. Define
@@
as the subset of $\mathcal{D}$ that contains messages in both categories. The first lemma below shows that most of the data contain both categories of messages with high probability.
@@
Under Assumption 1, if $P_{\overline{\pi}}(\mathcal{M}_{1})\in(0,1)$, $|\mathcal{Y}|\geq 2$, then there exists a constant $\gamma=\gamma(|\mathcal{Y}|)\in(0,1)$ such that for all $|\mathcal{D}|>0$, $\mathbb{P}(|\mathcal{I}|\leq|\mathcal{D}|(1-\gamma))\leq e^{-\frac{|\mathcal{D}|(1-\gamma)\gamma^{2}}{8}}$.
@@
Under Assumption 1, for each $(\mathcal{Y},y)\in\mathcal{D}$, the membership $\tau$ of each message $y^{\prime}\in\mathcal{Y}$ follows $\mathrm{Bernoulli}(P_{\overline{\pi}}(\mathcal{M}_{1}))$. Let $s=1-P_{\overline{\pi}}(\mathcal{M}_{1})^{|\mathcal{Y}|}-P_{\overline{\pi}}(\mathcal{M}_{2})^{|\mathcal{Y}|}<1$. Then $s$ equals the probability of sampling both categories of messages in each datum, and $|\mathcal{I}|\sim\mathrm{Binomial}(|\mathcal{D}|,s)$. By Chernoff’s bound,
$\mathbb{P}\left(|\mathcal{I}|\leq s\cdot|\mathcal{D}|s\right)\leq e^{-\frac{|\mathcal{D}|s(1-s)^{2}}{2}}.$
Letting $\gamma=1-s^{2}\in(0,1)$, we have $s(1-s)^{2}>(1-\gamma)\gamma^{2}/4$. Therefore,
$\mathbb{P}\left(|\mathcal{I}|\leq|\mathcal{D}|(1-\gamma)\right)\leq e^{-\frac{|\mathcal{D}|(1-\gamma)\gamma^{2}}{8}}.$ (20)